OpenStudy (biohazard9064):

Help Please how do I Answer How do you determine the intercepts from an equation or graph?

OpenStudy (lynfran):

the intercepts on a graph..it where the graph cuts the x and y axis respectively ...

OpenStudy (biohazard9064):

Im lost

OpenStudy (biohazard9064):

I understand what your saying but how to compleat it

OpenStudy (biohazard9064):

complete*

OpenStudy (biohazard9064):

I got it thanks any way

OpenStudy (lynfran):

to determine intercepts from an equation example y=x^2+x-2. to find the y intercept we let x=0 that is y=(0)^2+(0)-2 the y intercept here is y=-2 ..to find the x intercept we let y=0 that is 0=x^2+x-2 x=1 and x=-2

OpenStudy (lynfran):